OpenStudy (theedwardsfamily):

Solve. Express the answer in simplest form. 1/5+ w =7 1/3 w=

OpenStudy (anonymous):

\[\frac{ 1 }{5 } + w =7\frac{ 1 }{ 3 }\] Is this it?

OpenStudy (anonymous):

The first thing we can do is find the least common multiple of the denominators of the fractions so we can work with them.

OpenStudy (theedwardsfamily):

yes

OpenStudy (theedwardsfamily):

ok

OpenStudy (theedwardsfamily):

what r the lcm?

OpenStudy (anonymous):

If we write down multiples of 3 and 5, we need the smallest multiple that they share in common: 3: 3, 6, 9, 12, 15, 18, 21 5: 5, 10, 15, 20, 25, 30

OpenStudy (anonymous):

Which number from this list is the first and smallest multiple that both 3 and 5 share

OpenStudy (theedwardsfamily):

15

OpenStudy (anonymous):

Good job :), now we can convert the fractions so that they are over 15

OpenStudy (anonymous):

For \[\frac{ 1 }{ 5 }\] What do we have to multiply 5 by to get to 15?

OpenStudy (theedwardsfamily):

5*3

OpenStudy (anonymous):

So we can do: \[\frac{ 1* 3 }{ 5 * 3 }\]

OpenStudy (theedwardsfamily):

3/15

OpenStudy (anonymous):

Now we have \[\frac{ 3 }{ 15 } + w = 7\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

We need to convert 7 1/3 into 15ths now

OpenStudy (theedwardsfamily):

so what nxt?

OpenStudy (anonymous):

We could just convert 7/1 into 15ths, 1/3, and add them together

OpenStudy (anonymous):

So for 7/1 what do we have to multiply 1 by to get to 15

OpenStudy (theedwardsfamily):

15*1 right?

OpenStudy (anonymous):

Yep :), so we have to do: \[\frac{ 7 * 15 }{ 1 * 15 }\]

OpenStudy (theedwardsfamily):

105/15

OpenStudy (anonymous):

Now we have to add 105/15 to 1/3 in 15ths. So what does the denominator of 3 need to be multiplied by to get 15?

OpenStudy (anonymous):

3 * ? = 15

OpenStudy (theedwardsfamily):

is it 5? im confused a bit

OpenStudy (anonymous):

Yes, you did it :)

OpenStudy (anonymous):

Since 3 * 5 = 15 we can do this to get 1/3 as a fraction over 15: \[\frac{ 1*5 }{ 3*5 }\]

OpenStudy (theedwardsfamily):

5/15

OpenStudy (anonymous):

Yep :), so now I can write out what we have so far:

OpenStudy (theedwardsfamily):

ok

OpenStudy (anonymous):

\[\frac{ 3 }{ 15 } + w = \frac{ 105 }{ 15 }+\frac{ 5 }{ 15 }\]

OpenStudy (anonymous):

We could first simplify the right side of the equation

OpenStudy (anonymous):

By adding the two fractions

OpenStudy (anonymous):

\[\frac{ 105 }{ 15 }+\frac{ 5 }{ 15 } = \frac{ ? }{ 15 }\]

OpenStudy (anonymous):

We would simply add across

OpenStudy (theedwardsfamily):

110/15

OpenStudy (anonymous):

\[\frac{ 3 }{ 15 } + w = \frac{ 110 }{ 15 }\]

OpenStudy (anonymous):

What might we do to get w alone on one side of the equation

OpenStudy (theedwardsfamily):

oh man! gtg

OpenStudy (anonymous):

We can get rid of the\[\frac{ 3 }{ 15 } \] by subtracting from both sides

OpenStudy (anonymous):

\[w = \frac{ 107 }{ 15 }\]

OpenStudy (theedwardsfamily):

ok is my answer 2 2/15?

OpenStudy (anonymous):

It would be \[\frac{ 107 }{ 15 }\] As an improper fraction or: \[7\frac{ 2 }{ 15 }\] As a mixed number

OpenStudy (theedwardsfamily):

so 7 2/15 is the answer?

OpenStudy (anonymous):

I would use the mixed number, yep :)

OpenStudy (theedwardsfamily):

thanks!!!

OpenStudy (anonymous):

\[7\frac{ 2 }{ 15 }\]